Find two points on the line 6y+3x=5
Found 2 solutions by josgarithmetic, math_tutor2020:
Answer by josgarithmetic(39618)

You find them! Pick x value and find corresponding y value. Solve the given equation for y.

Answer by math_tutor2020(3817)


There are infinitely many points on this line since there are infinitely many numbers we could replace x with (or we could replace y with).

Let's say we plugged in x = 0.
Solve for y.
6y+3x = 5
6y+3(0) = 5
6y = 5
y = 5/6
When x = 0, it leads to y = 5/6
One point on this line is (0, 5/6) which is the y-intercept.
The y-intercept is where the line crosses the y axis.

Now let's plug in y = 0 and solve for x.
6y+3x = 5
6(0)+3x = 5
3x = 5
x = 5/3
The x-intercept is located at (5/3,0)
The x-intercept is where the line crosses the x axis.

Graph

Desmos and GeoGebra are two graphing options I recommend.
5/6 = 0.8333 approximately
5/3 = 1.6667 approximately

There isn't a rule you have to use zero. It just happens to be the easiest value to work with.
Feel free to try other values such as x = 1 or y = 5.